# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@
Abstract.
Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization. Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conducted the Systematic Mapping Study (SMS) on the literature published between January 2013 and July 2024. One hundred twenty one studies were selected and the key results are: (1) Four leading themes on containerization in multi-cloud environment are identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Ninety-eight patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four catalogs consisting of challenges and solutions related to security, automation, deployment, and monitoring were introduced. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions for containerization applications in multi-cloud environment.
Containerization, Multi-Cloud Environment, Systematic Mapping Study, Cloud Computing
journal: TOSEMccs: Software and its engineering Software development techniques
1. Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b). Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014). Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification (Docker, 2021). On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013). Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost-efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
@@ 1. Introduction
Figure 1. Context: Containerization in multi-cloud environment
@@
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022). These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018). Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
@@
Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design (Ahmad et al., 2025). The project brings together a consortium of four companies (Bittium [^1] https://www.bittium.com/, M-Files [^2] https://www.m-files.com/, Solita [^3] https://www.solita.fi/en/, and Vaadin [^4] https://vaadin.com/), each with distinct requirements for containerization, and is further supported by industry leaders Nokia [^5] https://www.nokia.com/ and TietoEVRY [^6] https://www.tietoevry.com/. This study fulfills an industry demand for optimized container strategies, addressing critical gaps in deployment consistency, security, and scalability that are necessary for practical applications in multi-cloud contexts. Despite the growing interest in containerization for multi-cloud environment, there remains a significant gap in systematically organized knowledge to address these challenges comprehensively. Our study not only fulfills an immediate industry demand but also aims to provide a structured view of containerization practices, addressing the needs of both practitioners and researchers in navigating this fragmented knowledge space. Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and gray literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive, secure solutions.
@@
Furthermore, this state of affairs of containerization in multi-cloud also poses challenges for the academic community. Researchers bear the obligation of navigating through this vast array of information to uncover the precise aspects they seek - be it a unique pattern, an unexplored challenge, or an innovative solution. Moreover, the distributed nature of this knowledge delays the formation of a cohesive and unified understanding of the subject matter. Recent studies (e.g., (Bordeleau et al., 2020) (Mihai et al., 2022) (Sarker, 2022)) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
@@
To bridge this knowledge gap and support the practical demands of the QLEAP project consortium, our study systematically identifies and categorizes the various facets of container utilization in multi-cloud environment. These facets include container roles, implementation strategies, architectural patterns, and quality attributes. In addition, we explore the challenges and solutions associated with automation, deployment, monitoring, and security, providing practitioners and researchers with actionable insights to tackle these critical issues effectively. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis. We adopted the following Goal-Question-Metric (GQM) framework (Caldiera and Rombach, 1994):
@@
  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS. This dataset includes detailed hierarchies of the developed catalogs, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
@@
2. Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007). Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process. Reliability and consistency are critical aspects of any empirical study. In this study, we followed established statistical approaches, including inter-rater agreement (IRA) computation, to ensure the reliability of our data coding and extraction processes. Additionally, we applied systematic quality assessment techniques to evaluate the studies included in our SMS.
Figure 2. Schematic representation of the research methodology implemented in this study
2.1. Research Questions
@@
Table 1. Research questions and their rationale
@@
2.2. Search String Composition
Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework (Schardt et al., 2007) (Ralph et al., 2020) for constructing our search string. However, given the complexity and breadth of our topic, applying PICO strictly required an extensive array of terms and logical operators, exceeding the query limits of databases like ScienceDirect. Therefore, we developed the search string for this study based on authors’ knowledge and iterative trial searches. This approach allowed us to tailor the search strategy to the specific goal of our study, ensuring comprehensive coverage within practical constraints.
2.3. Study Search and Selection Process
The study search and selection process for this study is divided into two phases. In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in (Waseem et al., 2024)). In the second phase, we applied the snowballing technique on primary studies that were selected from the first phase.
Table 2. Search string and targeted search area for databases
@@ 2.3. Study Search and Selection Process
2.3.1. Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and July 2024 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environment over the past decade. We followed the steps mentioned below to select the relevant studies.
Table 3. Inclusion and exclusion criteria for selecting primary studies in this SMS
@@ 2.3.1. Primary Search
Study Focus Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment. Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
@@
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 6,727 studies from eight databases. Following initial retrieval, we did not apply Cohen’s Kappa (Cohen, 1960) (Cohen, 1968) as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. Initially, we removed any duplicate studies obtained from different databases (e.g., IEEE Xplore, Scopus) by sorting them in ascending order. This resulted in the elimination of several thousand duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objective, as detailed in Table 3.
    To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1960) (Cohen, 1968) was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
    (1) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 (Sim and Wright, 2005). Specifically, R1 and R2 agreed on 360 studies being relevant (“YES”) and 20 studies being irrelevant (“NO”). Disagreements occurred in 20 studies, which were systematically resolved through discussions among all authors, leading to a final consensus. Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency (Keele et al., 2007; Chandler et al., 2019). Following this process, the number of studies was reduced to 1,047 for further analysis.
@@
    The outcomes of this assessment are depicted in Table 4.b. The analysis showed substantial agreement between R1 and R2, with 88 studies being unanimously considered relevant and 5 studies unanimously considered irrelevant. Disagreements occurred in 12 cases (7 studies where one researcher marked “relevant” while the other marked “irrelevant”, and 5 studies initially classified as “doubtful”). The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement (Sim and Wright, 2005; Keele et al., 2007).
    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant”, “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation. Specifically:
@@
    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines (Keele et al., 2007; Chandler et al., 2019). After resolving all disagreements through systematic discussions among the authors, a final consensus was reached for the subset. Consequently, upon completing this step, the number of studies was reduced to 346 for more detailed scrutiny.
@@
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 346 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly selected a subset of 35 studies (10% of the total 346 studies) for a Cohen’s Kappa analysis to measure the agreement between two researchers (R1 and R2). The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers (Sim and Wright, 2005). Specifically, R1 and R2 agreed on 30 studies as relevant (“YES”) and 2 studies as irrelevant (“NO”). Disagreements occurred in 3 studies (2 marked as “YES” by one researcher and “NO” by the other, and 1 marked as “NO” by both but flagged as uncertain). All disagreements were systematically resolved through discussions among the authors, leading to a final consensus. This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process (Chandler et al., 2019). At the end of this step, we obtained a total of 110 studies, which were deemed relevant for further analysis.
Table 4. Kappa Agreement across different screening steps.
(a) Title & Keyword Screening on 400 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 360 20
R2 No 10 10
Total 370 30
(b) Abstract-based Screening on 105 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 88 7
R2 No 5 5
Total 93 12
(c) Full-Text Screening on 35 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 30 3
R2 No 2 0
Total 32 3
2.3.2. Snowballing

### v3

@@
Abstract
Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conducted the Systematic Mapping Study (SMS) on the literature published between January 2013 and July 2024. One hundred twenty one studies were selected and the key results are: (1) Four leading themes on containerization in multi-cloud environment are identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Ninety-eight patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four catalogs consisting of challenges and solutions related to security, automation, deployment, and monitoring were introduced. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions for containerization applications in multi-cloud environment.
keywords:
[…]
1 Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years [1]. Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings [2]. Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification [3]. On the other hand, multi-cloud environment is the distribution of cloud assets [4]. Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost-efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
@@ 1 Introduction
Figure 1: Context: Containerization in multi-clouds
@@
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face [5]. These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds [6] [7], and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards [8]. Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively [9].
@@
Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design [10]. The project brings together a consortium of four companies (Bittium [^1] https://www.bittium.com/, M-Files [^2] https://www.m-files.com/, Solita [^3] https://www.solita.fi/en/, and Vaadin [^4] https://vaadin.com/), each with distinct requirements for containerization, and is further supported by industry leaders Nokia [^5] https://www.nokia.com/ and TietoEVRY [^6] https://www.tietoevry.com/. This study fulfills an industry demand for optimized container strategies, addressing critical gaps in deployment consistency, security, and scalability that are necessary for practical applications in multi-cloud contexts. Despite the growing interest in containerization for multi-cloud environment, there remains a significant gap in systematically organized knowledge to address these challenges comprehensively. Our study not only fulfills an immediate industry demand but also aims to provide a structured view of containerization practices, addressing the needs of both practitioners and researchers in navigating this fragmented knowledge space. Recent research (e.g., see Selected Studies sheet in [11]) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and gray literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive, secure solutions.
@@
Furthermore, this state of affairs of containerization in multi-cloud also poses challenges for the academic community. Researchers bear the obligation of navigating through this vast array of information to uncover the precise aspects they seek - be it a unique pattern, an unexplored challenge, or an innovative solution. Moreover, the distributed nature of this knowledge delays the formation of a cohesive and unified understanding of the subject matter. Recent studies (e.g., [12] [13] [14]) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
@@
To bridge this knowledge gap and support the practical demands of the QLEAP project consortium, our study systematically identifies and categorizes the various facets of container utilization in multi-cloud environment. These facets include container roles, implementation strategies, architectural patterns, and quality attributes. In addition, we explore the challenges and solutions associated with automation, deployment, monitoring, and security, providing practitioners and researchers with actionable insights to tackle these critical issues effectively. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis. We adopted the following Goal-Question-Metric (GQM) framework [15]:
@@
  • We have publicly released a dataset, which is available online [11], to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS. This dataset includes detailed hierarchies of the developed catalogs, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
@@
2 Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in [16] and augmenting them with SLR strategies [17]. Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process. Reliability and consistency are critical aspects of any empirical study. In this study, we followed established statistical approaches, including inter-rater agreement (IRA) computation, to ensure the reliability of our data coding and extraction processes. Additionally, we applied systematic quality assessment techniques to evaluate the studies included in our SMS.
Figure 2: Schematic representation of the research methodology implemented in this study
2.1 Research Questions
@@
Table 1: Research questions and their rationale
@@
2.2 Search String Composition
Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework [18] [19] for constructing our search string. However, given the complexity and breadth of our topic, applying PICO strictly required an extensive array of terms and logical operators, exceeding the query limits of databases like ScienceDirect. Therefore, we developed the search string for this study based on authors’ knowledge and iterative trial searches. This approach allowed us to tailor the search strategy to the specific goal of our study, ensuring comprehensive coverage within practical constraints.
2.3 Study Search and Selection Process
The study search and selection process for this study is divided into two phases. In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in [11]). In the second phase, we applied the snowballing technique on primary studies that were selected from the first phase.
Table 2: Search string and targeted search area for databases
@@ 2.3 Study Search and Selection Process
2.3.1 Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and July 2024 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged [20], marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environment over the past decade. We followed the steps mentioned below to select the relevant studies.
Table 3: Inclusion and exclusion criteria for selecting primary studies in this SMS
@@ 2.3.1 Primary Search
Study Focus Studies that explicitly address the intersection of containerization and multi-cloud environments. This includes studies that investigate the use of container technologies (e.g., Docker, Kubernetes) in the context of multi-cloud deployment, management, or architecture. Studies that focus solely on containerization without a multi-cloud context, or vice versa, are excluded. Studies that address only traditional cloud computing, single-cloud environments, or containerization without reference to multi-cloud contexts, and vice versa. Specifically, studies that examine container technologies without situating them in multi-cloud scenarios, or studies focused on multi-cloud architectures without discussing container-based solutions, are excluded.
@@
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 6,727 studies from eight databases. Following initial retrieval, we did not apply Cohen’s Kappa [21] [22] as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. To ensure alignment with our inclusion criteria, we manually reviewed study titles and abstracts to include only those that explicitly discuss containerization within multi-cloud contexts. Studies that focused solely on single-cloud containerization, general cloud orchestration without containers, or container use cases unrelated to multi-cloud environments were excluded—even if they matched keywords in the search string. Initially, we removed any duplicate studies obtained from different databases (e.g., IEEE Xplore, Scopus) by sorting them in ascending order. This resulted in the elimination of several thousand duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objective, as detailed in Table 3. To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis [21] [22] was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
    $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (1)
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 [23]. Specifically, R1 and R2 agreed on 360 studies being relevant (“YES”) and 20 studies being irrelevant (“NO”). Disagreements occurred in 20 studies, which were systematically resolved through discussions among all authors, leading to a final consensus. Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency [17, 24]. Following this process, the number of studies was reduced to 1,047 for further analysis.
@@
    The outcomes of this assessment are depicted in Table 4.b. The analysis showed substantial agreement between R1 and R2, with 88 studies being unanimously considered relevant and 5 studies unanimously considered irrelevant. Disagreements occurred in 12 cases (7 studies where one researcher marked “relevant” while the other marked “irrelevant”, and 5 studies initially classified as “doubtful”). The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement [23, 17].
    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant", “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation. Specifically:
@@
    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines [17, 24]. After resolving all disagreements through systematic discussions among the authors, a final consensus was reached for the subset. Consequently, upon completing this step, the number of studies was reduced to 346 for more detailed scrutiny.
@@
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 346 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly selected a subset of 35 studies (10% of the total 346 studies) for a Cohen’s Kappa analysis to measure the agreement between two researchers (R1 and R2). The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers [23]. Specifically, R1 and R2 agreed on 30 studies as relevant (“YES”) and 2 studies as irrelevant (“NO”). Disagreements occurred in 3 studies (2 marked as “YES” by one researcher and “NO” by the other, and 1 marked as “NO” by both but flagged as uncertain). All disagreements were systematically resolved through discussions among the authors, leading to a final consensus. This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process [24]. At the end of this step, we obtained a total of 110 studies, which were deemed relevant for further analysis.
Table 4: Kappa agreement across different screening steps
(a) Title & Keyword
Screening (400 Studies)
R1 Yes R1 No R2 Yes 360 20 R2 No 10 10 Total 370 30
(b) Abstract-Based
Screening (105 Studies)
R1 Yes R1 No R2 Yes 88 7 R2 No 5 5 Total 93 12
(c) Full-Text
Screening (35 Studies)
R1 Yes R1 No R2 Yes 30 3 R2 No 2 0 Total 32 3
2.3.2 Snowballing